Circular No. 5652

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bi IN ANONYMOUS GALAXY C. Pennypacker, S. Perlmutter, G. Goldhaber, A. Goobar, J. Desai,

A. Kim, M. Kim, R. Muller, H. Newberg, and I. Small, Lawrence Berkeley Laboratory and the Center for Particle Astrophysics, Berkeley; B. Boyle, R. McMahon, Institute of Astronomy, Cambridge; P. Bunclark, D. Carter, M. Irwin, and R. Terlevich, Royal Greenwich Observatory; R. Ellis, University of Durham; and W. Couch, University of New South Wales, report the detection of a supernova with the Isaac Newton 2.5-m telescope during a search for high-redshift supernovae. The supernova is at α = 16^{h}08^{m}28^{s}.40, δ = +39^{o}54′57′′.7 (equinox 1950.0), and is 1′′.5 east and 0′′.5 north of the core of the host galaxy. The supernova was discovered at magnitude R = 22 on an Apr. 21 image, and then confirmed on images from Apr. 24, 28, May 2, 7, 9, and 11. Two previous reference images from Mar. 25 and 27, and later follow-up images on June 10 (T. Small and J. Mould, Palomar 5-m reflector), July 6 (C. Crawford and R. Abraham, Isaac Newton telescope), and Aug. 21 and 22, all show no supernova to a limit of mag ∼ 24. The photometry points and upper limits indicate SN 1992bi to be a type-Ia supernova, with an implied redshift of ∼ 0.45. On Aug. 21, the spectrum of the host galaxy was observed at the William Herschel 4.2-m telescope by K. Glazebrook (Durham) and the redshift measured to be 0.457.

SUPERNOVAE 1992aa, 1992ab, 1992ax D. D. Balam, University of Victoria, reports his astrometric measure-

ments of supernovae (equinox 1950.0) from CCD images obtained by F. D. A. Hartwick, G. C. L. Aikman, and himself with the 1.85-m telescope of the Dominion Astrophysical Observatory: SN 1992aa in NGC 6464 (IAUC 5534), α = 17^{h}45^{m}13^{s}.30, δ = +60^{o}54′42′′.2; SN 1992ab in NGC 6389 (ibid.), α = 17^{h}30^{m}30^{s}.93, δ = +16^{o}25′49′′.6; SN 1992ax in anonymous galaxy (IAUC 5607), α = 22^{h}08^{m}24^{s}.27, δ = +44^{o}47′44′′.5 (host galaxy has end figures 23^{s}.98, 48′04′′.5).

NOVA SAGITTARII 1992 No. 3 Visual magnitude estimates: Oct. 18.45 UT, 8.1 (D. Seargent, The En-

trance, N.S.W.); 24.40, 8.6 (Seargent); 26.70, 9.1 (M. Tombelli, Montelupo, Italy); Nov. 4.72, 10.0 (A. Boattini, Florence, Italy).

1992 November 12 Daniel W. E. Green